# XMainframe: A Large Language Model for Mainframe Modernization

### v2

@@ 1 Introduction
  • Legacy Code Interpretation: Developers today encounter significant challenges when working with code written in outdated and legacy languages. These legacy systems often lack comprehensive documentation, making it difficult to interpret the original intent and functionality of the code. Furthermore, the original developers may no longer be available, leading to a substantial knowledge gap. CodeLLMs can serve as support tools capable of accurately summarizing or describing legacy code, helping developers understand and work with these systems more effectively. By generating detailed explanations and summaries, CodeLLMs bridge the knowledge gap and facilitate the interpretation of complex legacy code.
@@ 2.1 Code Large Language Models
There have been numerous Code-LLMs trained on massive datasets to achieve remarkable advancements across diverse coding tasks, including code generation Roziere et al. (2023); Touvron et al. (2023); Li et al. (2023); Jiang et al. (2024); Feng et al. (2020), code summarization Ahmed & Devanbu (2022); Lu et al. (2021); Gao et al. (2023); Su & McMillan (2024), program repair Xia & Zhang (2022); Wei et al. (2023); Xia et al. (2023), etc. Besides, LLMs showcase unforeseen capabilities that allow them to adapt to different domains without modifying their parameters by using discrete prompting. These prompts created by humans or by LLMs themselves involve providing instructions along with relevant contexts as input to refine the generation Luo et al. (2023); Wang et al. (2023a). Related to instruction tuning is chain-of-thought promoting, where models are encouraged to explain the reasoning when given a complex problem, thereby increasing the likelihood that their final answer is correct Wei et al. (2022). Recently, several studies proposed multi-agent collaborations, where each agent focuses on a unique task, such as code generation or task planning, to enhance the effectiveness of LLM-based code generation Chen et al. (2023); Qian et al. (2023); Huang et al. (2023).
@@ 2.2 LLMs for domain-specific tasks
Although general LLMs were trained to cover a broad range of topics, they were beaten by much smaller models using only domain-specific data on tasks within those domains Wu et al. (2023); Pal et al. (2024); Arefeen et al. (2024). Training specialized LLMs valued in various areas, such as finance Wu et al. (2023); Yang et al. (2023), law Cui et al. (2023), health Yang et al. (2022); Peng et al. (2023), IT operations Guo et al. (2023),… Therefore, it sheds light on the benefits and necessity of tailoring AI models to specific fields.
@@
Although understanding mainframe systems is crucial, it is rare to find an AI model to support this task. Granite Mishra et al. (2024) from IBM is the first model developed for this purpose. However, Granite has two limitations: it only maintains IBM’s Z-system and focuses primarily on documents rather than source code. As a result, it delivers low-quality performance on coding tasks for legacy systems, such as code completion or code summarization of the COBOL codebase. Mainframer from BloopAI Gordon-Hall (2024) is also one of the few models to support a coding task for a legacy COBOL system, achieving good performance on COBOL code completion. Nevertheless, it is trained only on a dataset specific to code completion, resulting in almost zero accuracy for other tasks like question and answering or code summarization. Unlike them, we aim to build a universal model that supports various tasks in this domain with high performance.
2.3 Benchmark for COBOL and mainframe systems.
@@
Code-related datasets have been introduced to verify empirical research across various programming languages and tackle challenges in multiple areas of software engineering Odena et al. (2021); Iyer et al. (2018); Chen et al. (2021); Nguyen et al. (2023). However, low-resource language such as COBOL has not received enough attention from scientific and academic researchers. Therefore, a big barrier exists to training LLMs for COBOL on a large scale. OpenCBS Lee et al. (2022) is one of the pioneering efforts leveraging public forums to create a COBOL set for defect detection. Another dataset, X-COBOL Ali et al. (2023), comprises 84 COBOL repositories collected from GitHub. Despite undergoing a data extraction pipeline, this dataset falls short of the intended quality as the authors relied on GitHub stars for filtering the repositories, which is not a sufficient metric for determining the quality of a repository Borges & Valente (2018). More recently, BloopAI announced the COBOLEval Gordon-Hall (2024), a benchmark designed to evaluate legacy code completion tasks. It consists of 146 coding problems converted into COBOL from HumanEval benchmark Chen et al. (2021), which is a Python dataset. Nevertheless, it is unrealistic since COBOL is primarily used for business and finance systems, not for solving programming challenges. To the best of our knowledge, there is no dataset that covers diverse tasks related to COBOL language and legacy systems.
@@ 3.1 Dataset for Pretraining
Figure 1: Data Augmentation Pipeline
We initially retrieved all GitHub repositories containing COBOL and Mainframe system code, amassing approximately 4GB of data. To ensure high-quality training samples, we removed overly short repositories and files, eliminated alphanumeric character fractions, binary data, JSON, XML data, and node modules, resulting in 40,960 COBOL files. We further refined our dataset using MinHash and Locality Sensitive Hashing (LSH) to detect and remove near-duplicates [citation]. This process involved document shingling and fingerprinting, using locality-sensitive hashing to group similar documents, detecting actual duplicates, and removing them. The final COBOL dataset consists of 33,561 files, encompassing 228 million tokens in 8 million Lines of Code (LoCs). For Mainframe documents, we extracted data from public books and websites related to Mainframe and COBOL, ensuring minimal noise and maximum data cleanliness. We extracted main content from HTML pages and eliminated unnecessary parts using specific tags, IDs, and keywords, resulting in 14,274 documents containing approximately 8 million tokens. Copy
@@ 3.2 Dataset for Model Instruct
In the initial stage, 300 seed data instances about Mainframe and COBOL are gathered and annotated by our domain experts in the QA formats Zhang et al. (2023). They include knowledge-based question answering, deployment, syntax, COBOL code summarization, and other aspects. To cover various practical scenarios, we design long and short versions with different styles of instruction prompts for each type of question, assisting with our large language model’s supervised instruction-tuning process. Then, we use OpenAI GPT-4-turbo to generate more than 200 sub-topics within the Mainframe and COBOL fields. This step is designed to ensure the generated content is firmly rooted in our specific domain. All of them serve as the foundation for data augmentation, enhancing the scale and diversity of our dataset.
@@ 5.3.1 Multiple Choice Question
Table 2: Accuracy results on the multiple-choice question task.

### v3

@@ 1 Introduction
  • Legacy Code Interpretation: Developers today face significant challenges when working with code written in outdated and legacy languages. These legacy systems often lack comprehensive documentation, making it difficult to understand the original intent and functionality of the code. Additionally, the original developers may no longer be available, creating a substantial knowledge gap. To address this, AI systems that assist developers in interpreting legacy codebases must be capable of understanding code at a repository-level scale Phan et al. (2024); Zhang et al. (2023a); Liu et al. (2024). Such tools can provide accurate summaries and descriptions of legacy code, enabling developers to work with these systems more effectively. By generating detailed explanations and summaries, CodeLLMs help bridge the knowledge gap, facilitating the interpretation and maintenance of complex legacy code.
@@ 2.1 Code Large Language Models
Numerous Code-LLMs have been trained on massive datasets, leading to significant advancements across various coding tasks, including code generation Roziere et al. (2023); Touvron et al. (2023); Li et al. (2023); Jiang et al. (2024); Feng et al. (2020), code summarization Ahmed & Devanbu (2022); Lu et al. (2021); Gao et al. (2023); Su & McMillan (2024); To et al. (2023); Bui & Jiang (2018); Nguyen et al. (2022), and program repair Xia & Zhang (2022); Wei et al. (2023); Xia et al. (2023); Bui et al. (2022). These models have also demonstrated unexpected capabilities, such as adapting to different domains through discrete prompting, without requiring parameter modifications. Human-crafted or LLM-generated prompts, which include instructions and relevant context, are used to refine the generation process Luo et al. (2023); Wang et al. (2023a). Related to instruction tuning is chain-of-thought prompting, where models are encouraged to explain their reasoning when faced with complex problems, increasing the likelihood of correct answers Wei et al. (2022). Recently, several studies have explored multi-agent collaborations, where each agent specializes in a unique task—such as code generation or task planning—to enhance the effectiveness of LLM-based code generation Chen et al. (2023); Qian et al. (2023); Huang et al. (2023).
@@ 2.2 LLMs for Domain-Specific Tasks
While general LLMs are trained to cover a wide range of topics, they are often outperformed by smaller models trained exclusively on domain-specific data in tasks within those domains Wu et al. (2023); Pal et al. (2024); Arefeen et al. (2024). This has led to the development of specialized LLMs in various areas, such as finance Wu et al. (2023); Yang et al. (2023), law Cui et al. (2023), health Yang et al. (2022); Peng et al. (2023), and IT operations Guo et al. (2023). The success of these models underscores the benefits and necessity of tailoring AI models to specific fields.
@@
In the context of mainframe systems, which are critical yet underrepresented in AI research, there are very few AI models designed to support tasks in this domain. Granite Mishra et al. (2024) from IBM is the first model developed for this purpose. However, Granite has limitations: it only supports IBM’s Z-system and focuses primarily on documents rather than source code, resulting in suboptimal performance on coding tasks for legacy systems, such as code completion or code summarization of COBOL codebases. Another model, Mainframer from BloopAI Gordon-Hall (2024), is one of the few models designed to support coding tasks for legacy COBOL systems, achieving good performance in COBOL code completion. However, it is trained solely on a dataset specific to code completion, rendering it nearly ineffective for other tasks like question answering or code summarization. In contrast, our goal is to build a universal model that excels across various tasks in this domain, delivering high performance consistently.
2.3 Benchmark for COBOL and Mainframe Systems
@@
Code-related datasets have been developed to facilitate empirical research across various programming languages and address challenges in multiple areas of software engineering Odena et al. (2021); Iyer et al. (2018); Chen et al. (2021); Nguyen et al. (2023). However, low-resource languages like COBOL have received limited attention from the scientific and academic communities, creating a significant barrier to training LLMs for COBOL on a large scale. OpenCBS Lee et al. (2022) is one of the pioneering efforts in this space, leveraging public forums to create a COBOL dataset for defect detection. Another dataset, X-COBOL Ali et al. (2023), consists of 84 COBOL repositories collected from GitHub. Despite undergoing a data extraction pipeline, this dataset falls short in quality because the authors relied on GitHub stars for filtering repositories, which is not a reliable metric for determining repository quality Borges & Valente (2018). More recently, BloopAI announced COBOLEval Gordon-Hall (2024), a benchmark designed to evaluate legacy code completion tasks. It consists of 146 coding problems converted into COBOL from the HumanEval benchmark Chen et al. (2021), originally a Python dataset. However, this approach is unrealistic, as COBOL is primarily used in business and finance systems, not for solving general programming challenges. To the best of our knowledge, there is no dataset that comprehensively covers diverse tasks related to the COBOL language and legacy systems.
@@ 3.1 Dataset for Pretraining
Figure 1: Data Augmentation Pipeline.
We initially retrieved all GitHub repositories containing COBOL and Mainframe system code, amassing approximately 4GB of data. To ensure high-quality training samples, we removed overly short repositories and files, eliminated alphanumeric character fractions, binary data, JSON, XML data, and node modules, resulting in 40,960 COBOL files. We further refined our dataset using MinHash and Locality Sensitive Hashing (LSH) to detect and remove near-duplicates [citation]. This process involved document shingling and fingerprinting, using locality-sensitive hashing to group similar documents, detecting actual duplicates, and removing them. The final COBOL dataset consists of 33,561 files, encompassing 228 million tokens in 8 million Lines of Code (LoCs). For Mainframe documents, we extracted data from public books and websites related to Mainframe and COBOL, ensuring minimal noise and maximum data cleanliness. We extracted main content from HTML pages and eliminated unnecessary parts using specific tags, IDs, and keywords, resulting in 14,274 documents containing approximately 8 million tokens.
@@ 3.2 Dataset for Model Instruct
In the initial stage, 300 seed data instances about Mainframe and COBOL are gathered and annotated by our domain experts in the QA formats Zhang et al. (2023b). They include knowledge-based question answering, deployment, syntax, COBOL code summarization, and other aspects. To cover various practical scenarios, we design long and short versions with different styles of instruction prompts for each type of question, assisting with our large language model’s supervised instruction-tuning process. Then, we use OpenAI GPT-4-turbo to generate more than 200 sub-topics within the Mainframe and COBOL fields. This step is designed to ensure the generated content is firmly rooted in our specific domain. All of them serve as the foundation for data augmentation, enhancing the scale and diversity of our dataset.
@@ 5.3.1 Multiple Choice Question
Table 2: Results on Multiple-Choice Question.